OpenStudy (appleduardo):

whats the integral for x^4 / (1-x) ??

OpenStudy (klimenkov):

Try to get some common fractions from this integrand.

hartnn (hartnn):

i would have written (x^4-1)+1 as numerator then separate the numerator,-[ (x^4-1)/(x-1) +1/(x-1)] 2nd term is easy to integrate. for 1st term, you can cancel out common factor of (x-1) from numerator and denominator.

hartnn (hartnn):

if you know how to factorize x^4-1

OpenStudy (klimenkov):

Hint:\[x^4=x^4-1+1=(x^2-1)(x^2+1)+1=(x-1)(x+1)(x^2+1)-1.\]

hartnn (hartnn):

^

OpenStudy (anonymous):

or make a substitution as u=1-x